# New GPU developments in the Madgraph CUDACPP plugin: kernel splitting, helicity streams, cuBLAS color sums

### v2

@@ Abstract
The first production release of the CUDACPP plugin for the Madgraph5_aMC@NLO generator, which speeds up matrix element (ME) calculations for leading-order (LO) processes using a data parallel approach on vector CPUs and GPUs, was delivered in October 2024. This has been described in previous publications by the team behind that effort. In this paper, I describe my work on some additional developments and optimizations of CUDACPP, mainly but not exclusively for GPUs. The new approach, which represents a major restructuring of the CUDACPP computational engine, mainly consists in splitting the calculation of the ME, which was previously performed using a single large GPU kernel, into many smaller kernels. A first batch of changes, involving the move to separate “helicity streams” and the optional offloading of QCD color sums to BLAS, also via GPU tensor cores, was recently merged into a new production release of CUDACPP, in collaboration with my colleagues. Since then, I have completed a second batch of changes, which I now also consider ready to be merged in production, involving the possibility to split the calculation into groups of Feynman diagrams in separate source code files. This new feature makes it possible to compute QCD matrix elements for physics processes with a larger number of final state gluons: in particular, I present the first performance results from CUDACPP for the $2\!\rightarrow\!6$ process $gg\!\rightarrow\!t\bar{t}gggg$ on CPUs and GPUs and the $2\!\rightarrow\!7$ process $gg\!\rightarrow\!t\bar{t}ggggg$ on CPUs, which involve over 15k and 230k Feynman diagrams, respectively. I also take this opportunity to describe in detail some features of the CUDACPP software which had not yet been documented, both in the GPU and vector CPU implementations.
@@ 1 Introduction
In this context, a recent progress has been the port and optimization of the MG5aMC software for data parallel processing on graphical processing units (GPUs) and on CPUs with SIMD (Single Instruction Multiple Data) vector instructions. After a five-year development programme, which was initially facilitated by the activities of the HEP Software Foundation (HSF) event generator working group bib:csbs , and which regularly reported its progress through conference proceedings bib:vchep2021 ; bib:ichep2022 ; bib:acat2022 ; bib:chep2023 , this project delivered its first production release of the MG5aMC code with GPU and SIMD support for QCD leading-order (LO) processes in October 2024. This has been reported at the CHEP2024 conference bib:chep2024 and in a more recent journal submission bib:zw2025 . MG5aMC is a code-generating framework, largely written in Python, which allows users to generate code, by default using Fortran, for any physics process of their choice. For complex processes, the computational bottleneck of the MG5aMC workflow is the calculation of the matrix element (ME) for each phase space point from the momenta of the initial and final state particles in that event. The main outcome of this project has been a new code-generating plugin, which was named “CUDACPP” because the generated physics code uses C++ instead of Fortran to execute the calculation of MEs on vector CPUs, and also includes CUDA extensions to run the code on NVidia GPUs (as well as HIP extensions for AMD GPUs). The new code speeds up the ME computation by performing it for many events at the same time using a data parallel approach bib:hsf2020 ; bib:lhcc2020 . To achieve this, work was also needed in the process-agnostic core MG5aMC event processing algorithm for LO processes, MadEvent, which is written in Fortran and with which the process-specific C++ code generated by CUDACPP is interfaced and linked: the main change is the move from a sequential single-event API to a parallel multi-event API, and the refactoring of some functions to make them stateless and re-entrant. This effort has been the result of a collaboration involving many existing and new contributors of the MG5aMC project, as detailed in the references cited above bib:vchep2021 ; bib:ichep2022 ; bib:acat2022 ; bib:chep2023 ; bib:chep2024 ; bib:zw2025 .
@@
In this paper, I describe my work on many additional enhancements and optimizations of CUDACPP, mainly but not exclusively targeting GPUs. These are developments for which the initial brainstorming with some of my colleagues from the development team, notably Olivier Mattelaer who prototyped with me the first concrete changes to the code, took place during the CSCS GPU Hackathon that we attended in Lugano in September 2022 bib:hack2022 . I then invested further effort in this area in February and especially in October 2024 bib:ks2024 after the first CUDACPP release, but had to pause this work again due to other constraints. In September 2025, I was finally able to resume this R&D activity, which I now consider essentially complete and which I am documenting in detail in this paper in view of its possible integration in new production versions of the CUDACPP software. A first batch of changes, which I had described in the preprint of a previous version of this paper bib:preprintV1 , has already been merged into a new production release of CUDACPP bib:cudacpp364 , in collaboration with my colleagues. Since then, I have completed a second batch of changes, which were still work in progress at the time of Ref. bib:preprintV1 , but which I now also consider ready to be merged in production.
@@ 1.1 Kernel splitting
The main idea behind all of these developments can be referred to as “kernel splitting”. In short, this consists in replacing the large monolithic kernel previously used by CUDACPP for the calculation of MEs from particle momenta, named sigmaKin, by many smaller kernels, executed either in parallel or sequentially, to achieve a more efficient and scalable use of the GPU. The aim is not only that of possibly achieving a higher peak throughput (in terms of MEs computed per second) with a large GPU grid, i.e. with a large number of events processed in parallel on the GPU during one offloading cycle, but also that of achieving a faster increase of throughput already with smaller GPU grids. This was the main focus of my first batch of changes, which have since been released. An important practical difference between the traditional MG5aMC workflow on CPUs and that using GPUs, in fact, is that the former was designed to handle iterations involving hundreds of events at a time, while the latter typically needs many thousands of events to be efficient: in this context, achieving higher throughputs with smaller grids would provide one way to use more manageable, but still reasonably efficient, event generation jobs on GPUs. An additional benefit of kernel splitting, which eventually became the main focus of my second batch of changes, is that smaller kernels are much more manageable by humans and compilers alike (both for CPUs and GPUs), and make it possible to achieve complex calculations that are instead very difficult if not impossible to perform with large monolithic kernels, such as the calculations of QCD matrix elements for physics processes with a large number of final state gluons. In this paper, in particular, I present the first performance results from CUDACPP for the $2\!\rightarrow\!6$ process $gg\!\rightarrow\!t\bar{t}gggg$ on CPUs and GPUs and for the $2\!\rightarrow\!7$ process $gg\!\rightarrow\!t\bar{t}ggggg$ on CPUs, which involve over 15k and over 230k Feynman diagrams, respectively.
@@
color sum: one kernel per helicity OR one BLAS per helicity
@@
color sum: one kernel per helicity OR one single BLAS for all helicities
@@
Figure 1: Schematic representation of the architectural evolution of matrix element calculations in the MG5aMC CUDACPP plugin, in the context of the kernel splitting developments presented in this paper. The steps above the blue dotted line, up to ihel3 and ihel4 included, have been presented in a previous preprint bib:preprintV1 of this paper, while those below it are described here for the first time. The steps in the green boxes, up to ihel3p1 included, have been merged upstream and included in production releases of CUDACPP; the steps in the red boxes, conversely, from ihel4 to ihel6p2 included, are included in a pull request that I recommend for merging upstream. More details about each step are provided in the text and in other figures.
@@
Schematically, in this paper I describe my work on six sets of kernel splitting developments: (1) helicity streams; (2) color sum splitting into separate GPU kernels; (3) color sum host refactoring with optional BLAS offloading; (4) Feynman diagram splitting, with one separate GPU kernel for each diagram; (5) Feynman diagram splitting, with one GPU kernel for each group of diagrams, and separate source code files; (6) Feynman diagram splitting, with a single GPU kernel processing all diagram groups as GPU device functions. A brief description of these steps and of the connection between them is shown in Fig. 1. The performance of each set of developments was tested for different physics processes, software configurations and hardware architectures. While these developments were initially aimed at improving throughputs on GPUs, the large code refactoring that they imply was also propagated to the vectorised C++ code for SIMD CPUs, as code generated by the CUDACPP plugin is strictly the same for CPUs and GPUs, and the distinction between the two cases is only done at build time through #ifdef directives. I therefore analysed not only CUDA and HIP performance on NVidia and AMD GPUs, but also C++ performance on vector CPUs using different SIMD configurations. The possibility to split the ME calculation into groups of Feynman diagrams in different files, in particular, was the main enabling factor for being able to compute complex $2\!\rightarrow\!7$ physics processes on CPUs.
@@
Considering the results of all my tests, my recommendation is to include the full set of these enhancements in new production releases of CUDACPP. This differs from the situation that I had described in my previous preprint bib:preprintV1 , where the diagram splitting work had not yet been completed. A first batch of changes (green boxes in Fig. 1) has already been merged into a new production release, in collaboration with my colleagues: this includes the first three steps of this work, notably helicity streams, the separation of color sums and Feynman diagrams, and the optional BLAS offloading of color sums (which was significantly improved with respect to that described in Ref. bib:preprintV1 ). The second batch of developments (red boxes in Fig. 1), notably including the splitting of Feynman diagrams into different groups and source code files, is in my opinion now also ready to be merged upstream. I note, in particular, that the performance of the code in its default configuration would not be negatively impacted by these changes: many of the features that I have added, in fact, are optional and are only enabled with specific flags at generation time, build time or run time (this includes the BLAS offloading of color sums, the ME splitting into groups of Feynman diagrams, their orchestration via CUDA graphs and the choice to treat diagram groups as separate kernels or as device functions within a single kernel).
@@ 1.2 Outline of this paper
As these new developments imply a significant refactoring of the existing CUDACPP codebase, and rely on internal features that have not yet been documented, I also found it useful to start this article with a brief review of some architectural aspects of the software, focusing on those most relevant to the kernel splitting developments, but also covering some details of the SIMD implementation. In this context, I should mention that the internal design of the CUDACPP plugin, which was largely my own work, is far from perfect and the code could certainly benefit from an extensive cleanup. In particular, over time I have added various switches to support alternative implementations of some software components, as I have done again in many of the new developments described in this paper: I believe that this has been very useful, and in some cases essential, to try out different approaches and eventually converge on well-defined solutions to address specific problems, but I am also aware that this results in code that might be more difficult to read, and where some of this flexibility may be regarded as an unnecessary complication. I hope that this article may help clarify the rationale behind some of these design choices and provide some guidance for possible future developments and/or cleanups. I stress, in any case, that the opinions that I express in this paper are only my own, and may differ from those of some of my colleagues.
The outline of this paper is the following. In Sec. 2, I briefly review some aspects of the CUDACPP software architecture design, focusing on those relevant to GPU kernel splitting, but also covering some details of the SIMD implementation for vector CPUs. In Sec. 3, I describe in detail the software design and implementation of my first batch of kernel splitting enhancements, and I provide the results and an analysis of various performance tests: this includes and partially extends the material that I had covered in my previous preprint bib:preprintV1 , with the exception of Feynman diagram splitting. In Sec. 4, I describe the various options I introduced to split the calculation of Feynman diagrams into different groups of diagrams and/or different kernels, and different source code files; I also give the first CUDACPP performance results for physics processes with large numbers of final state gluons, which could be computed thanks to these new features. In Sec. 5, I finally give my conclusions and an an outlook for this work.
@@ 2.1 Three software areas and development phases
    ME engines: the port to GPUs and SIMD CPUs, using CUDA/HIP and vectorised C++, of the computational engines to calculate MEs from particle momenta, and their backport to the code-generating Python framework, for different physics processes.
@@
More specifically, in all of the work that I present in this paper, I only developed code using the standalone application check.exe, where all three main components of an event generation application (random numbers, phase space sampling and ME calculation) are implemented in C++ and CUDA using a data parallel approach. This is because check.exe makes it possible to focus on the optimization of the ME engine without being constrained by the MadEvent Fortran infrastructure, where the legacy phase space sampling and other sequential non-ME components slow down the whole workflow. I routinely built check.exe using the CUDACPP makefile cudacpp.mk, without even building MadEvent or any other Fortran code (which proceeds through a separate makefile that internally delegates the build of the CUDACPP library to cudacpp.mk).
@@
Table 1: Pseudo-code of the ME calculation in CUDACPP before any kernel splitting changes (“ihel0”). Function main represents the check.exe standalone application. Note that sigmaKin is the only GPU kernel (__global__ function) in this workflow. Only a simplified GPU version of the code is shown, using scalar fptype floating point and cxtype complex number types: the actual code that supports both GPUs and SIMD CPUs uses fptype_sv and cxtype_sv data types inside calculate_wavefunction, making it possible to use the same formal code for scalar data and for SIMD vectors (see Ref. bib:vchep2021 for details). The CPU branch of the code for main, encapsulated by an #ifdef and not shown here, uses the Host versions of various classes instead of the Device versions. The pseudo-code for the color sum schematically indicates that this proceeds via a quadratic form using a color matrix, and that color amplitudes are converted from fptype/cxtype floating point precision to a potentially different fptype2/cxtype2 precision (this enables the “mixed” precision mode, where the former is based on double and the latter on float).

### v3

@@ Abstract
The first production release of the CUDACPP plugin for the Madgraph5_aMC@NLO generator, which speeds up matrix element (ME) calculations for leading-order (LO) processes using a data parallel approach on vector CPUs and GPUs, was delivered in October 2024. This was described in previous publications by the team behind that effort. In this paper, I describe my work on some additional developments and optimizations of CUDACPP, mainly but not exclusively for GPUs. The new approach, which represents a major restructuring of the CUDACPP computational engine, primarily consists in splitting the ME calculation, previously performed using a single large GPU kernel, into many smaller kernels. A first batch of changes, involving the move to separate “helicity streams” and the optional offloading of QCD color sums to BLAS, was recently merged into a new CUDACPP release, in collaboration with my colleagues. Since then, I have completed a second batch of changes, involving the possibility to split the calculation into groups of Feynman diagrams in separate source code files. This new feature makes it possible to compute QCD matrix elements for physics processes with a larger number of final state gluons: in particular, I present the first performance results from CUDACPP for the $2\!\rightarrow\!6$ process $gg\!\rightarrow\!t\bar{t}gggg$ on CPUs and GPUs and the $2\!\rightarrow\!7$ process $gg\!\rightarrow\!t\bar{t}ggggg$ on CPUs, which involve over 15k and 230k Feynman diagrams, respectively. I also take this opportunity to describe in detail some previously undocumented features of the CUDACPP software, both in the GPU and vector CPU implementations.
@@ 1 Introduction
In this context, a recent progress has been the port and optimization of the MG5aMC software for data parallel processing on graphical processing units (GPUs) and on CPUs with SIMD (Single Instruction Multiple Data) vector instructions. After a five-year development programme, which was initially facilitated by the activities of the HEP Software Foundation (HSF) event generator working group bib:csbs , and which regularly reported its progress through conference proceedings bib:vchep2021 ; bib:ichep2022 ; bib:acat2022 ; bib:chep2023 , this project delivered its first production release of the MG5aMC code with GPU and SIMD support for leading-order (LO) processes in October 2024. This has been reported at the CHEP2024 conference bib:chep2024 and in a more recent journal submission bib:zw2025 . MG5aMC is a code-generating framework, largely written in Python, which allows users to generate code, by default using Fortran, for any physics process of their choice. For complex processes, the computational bottleneck of the MG5aMC workflow is the calculation of the matrix element (ME) for each phase space point from the momenta of the initial and final state particles in that event. The main outcome of this project has been a new code-generating plugin, which was named “CUDACPP” because the generated physics code uses C++ instead of Fortran to execute the calculation of MEs on vector CPUs, and also includes CUDA extensions to run the code on NVidia GPUs (as well as HIP extensions for AMD GPUs). The new code speeds up the ME computation by performing it for many events at the same time using a data parallel approach bib:hsf2020 ; bib:lhcc2020 . To achieve this, work was also needed in the process-agnostic core MG5aMC event processing algorithm for LO processes, MadEvent, which is written in Fortran and with which the process-specific C++ code generated by CUDACPP is interfaced and linked: the main change is the move from a sequential single-event API to a parallel multi-event API, and the refactoring of some functions to make them stateless and re-entrant bib:ecfa2023 . This effort has been the result of a collaboration involving many existing and new contributors of the MG5aMC project, as detailed in the references cited above bib:vchep2021 ; bib:ichep2022 ; bib:acat2022 ; bib:chep2023 ; bib:chep2024 ; bib:zw2025 .
@@
In this paper, I describe my work on many additional enhancements and optimizations of CUDACPP, mainly but not exclusively targeting GPUs. These are developments for which the initial brainstorming with some of my colleagues from the development team, notably Olivier Mattelaer who prototyped with me the first concrete changes to the code, took place during the CSCS GPU Hackathon that we attended in Lugano in September 2022 bib:hack2022 . I then invested further effort in this area in February 2024 and especially October 2024 bib:ks2024 after the first CUDACPP release, but had to pause this work again due to other constraints. In September 2025, I was finally able to resume this R&D activity, which I now consider essentially complete and which I am documenting in detail in this paper in view of its possible integration in new production versions of the CUDACPP software. A first batch of changes, which I had described in the preprint of a previous version of this paper bib:preprintV1 , has already been merged into a new production release of CUDACPP bib:cudacpp364 , in collaboration with my colleagues. Since then, I have completed a second batch of changes, which were still work in progress at the time of Ref. bib:preprintV1 , but which I now also consider ready to be merged in production.
@@ 1.1 Kernel splitting
The main idea behind these new developments can be referred to as “kernel splitting”. In short, this consists in replacing the large monolithic kernel previously used by CUDACPP for the calculation of MEs from particle momenta, named sigmaKin, by many smaller kernels, executed either in parallel or sequentially, to achieve a more efficient and scalable use of the GPU. The aim is not only that of possibly achieving a higher peak throughput (in terms of MEs computed per second) with a large GPU grid, i.e. with a large number of events processed in parallel on the GPU during one offloading cycle, but also that of achieving a faster increase of throughput already with smaller GPU grids. This was the main focus of my first batch of changes, which have since been released. An important practical difference between the traditional MG5aMC workflow on CPUs and that using GPUs, in fact, is that the former was designed to handle iterations involving hundreds of events at a time, while the latter typically needs many thousands of events to be efficient: in this context, achieving higher throughputs with smaller grids would provide one way to use more manageable, but still reasonably efficient, event generation jobs on GPUs. An additional benefit of kernel splitting, which eventually became the main focus of my second batch of changes, is that smaller kernels are much more manageable by humans and compilers alike (both for CPUs and GPUs), and make it possible to achieve complex calculations that are instead very difficult if not impossible to perform with large monolithic kernels, such as the calculations of QCD matrix elements for physics processes with a large number of final state gluons. In this paper, in particular, I present the first performance results from CUDACPP for the $2\!\rightarrow\!6$ process $gg\!\rightarrow\!t\bar{t}gggg$ on CPUs and GPUs and for the $2\!\rightarrow\!7$ process $gg\!\rightarrow\!t\bar{t}ggggg$ on CPUs, which involve over 15k and over 230k Feynman diagrams, respectively.
@@
GPU color sum: one kernel per helicity OR one BLAS per helicity
@@
GPU color sum: one kernel per helicity OR one single BLAS for all helicities
csm
CPU color sum: FPTYPE=m SIMD optimizations
@@
Figure 1: Schematic representation of the architectural evolution of matrix element calculations in the MG5aMC CUDACPP plugin, in the context of the kernel splitting developments presented in this paper. The steps above the blue dotted line, up to ihel3 and ihel4 included, have been presented in the first preprint bib:preprintV1 of this paper, while those below it are only described here. The steps in the green boxes (up to ihel3p1 included) have been merged upstream and included in production releases of CUDACPP; the steps in the red boxes (from ihel4 to ihel6p2 included) and the step in the orange box (csm), conversely, are included in two pull requests that I recommend for merging upstream. More details about each step are provided in the text and in other figures.
@@
Schematically, in this paper I describe my work on six sets of kernel splitting developments: (1) helicity streams; (2) color sum splitting into separate GPU kernels; (3) color sum host refactoring with optional BLAS offloading; (4) Feynman diagram splitting, with one separate GPU kernel for each diagram; (5) Feynman diagram splitting, with one GPU kernel for each group of diagrams, and separate source code files; (6) Feynman diagram splitting, with a single GPU kernel processing all diagram groups as GPU device functions. A brief description of these steps and of the connection between them is shown in Fig. 1. The performance of each set of developments was tested for different physics processes, software configurations and hardware architectures. While these developments were initially aimed at improving throughputs on GPUs, the large code refactoring that they imply was also propagated to the vectorized C++ code for SIMD CPUs, as code generated by the CUDACPP plugin is strictly the same for CPUs and GPUs, and the distinction between the two cases is only done at build time through #ifdef directives. I therefore analysed not only CUDA and HIP performance on NVidia and AMD GPUs, but also C++ performance on vector CPUs using different SIMD configurations. The possibility to split the ME calculation into groups of Feynman diagrams in different files, in particular, was the main enabling factor for being able to compute complex $2\!\rightarrow\!7$ physics processes on CPUs. In turn, this also pushed me to further analyse and optimize a few specific details of the SIMD implementation for vector CPUs, notably the QCD color sum.
@@
Considering the results of all my tests, my recommendation is to include the full set of these enhancements in new production releases of CUDACPP. This differs from the situation that I had described in my first preprint bib:preprintV1 , where the diagram splitting work had not yet been completed. A first batch of changes (green boxes in Fig. 1) has already been merged into a new production release, in collaboration with my colleagues: this includes the first three steps of this work, notably helicity streams, the separation of color sums and Feynman diagrams, and the optional BLAS offloading of color sums (which was significantly improved with respect to that described in Ref. bib:preprintV1 ). The second batch of developments (red and orange boxes in Fig. 1), notably including the splitting of Feynman diagrams into different groups and source code files, is in my opinion now also ready to be merged upstream. I note, in particular, that the performance of the code in its default configuration would not be negatively impacted by these changes: many of the features that I have added, in fact, are optional and are only enabled with specific flags at generation time, build time or run time (this includes the BLAS offloading of color sums, the ME splitting into groups of Feynman diagrams, their orchestration via CUDA graphs and the choice to treat diagram groups as separate kernels or as device functions within a single kernel).
@@ 1.2 Outline of this paper
As these new developments imply a significant refactoring of the existing CUDACPP codebase, and rely on internal features that have not yet been documented, I also found it useful to start this article with a brief review of some architectural aspects of the software, focusing on those most relevant to the GPU kernel splitting developments, but also covering some details of the SIMD implementation. In this context, I should mention that the internal design of the CUDACPP plugin, which was largely my own work, is far from perfect and the code could certainly benefit from an extensive cleanup. In particular, over time I have added various switches to support alternative implementations of some software components, as I have done again in many of the new developments described in this paper: I believe that this has been very useful, and in some cases essential, to try out different approaches and eventually converge on well-defined solutions to address specific problems, but I am also aware that this results in code that might be more difficult to read, and where some of this flexibility may be regarded as an unnecessary complication. I hope that this article may help clarify the rationale behind some of these design choices and provide some guidance for possible future developments and/or cleanups. I stress, in any case, that the opinions that I express in this paper are only my own, and may differ from those of some of my colleagues.
The outline of this paper is the following. In Sec. 2, I briefly review some aspects of the CUDACPP software architecture design, focusing on those relevant to GPU kernel splitting, but also covering some details of the SIMD implementation for vector CPUs. In Sec. 3, I describe in detail the software design and implementation of my first batch of kernel splitting enhancements, and I give the results and an analysis of various performance tests: this includes and partially extends the material that I had covered in my first preprint bib:preprintV1 , with the exception of Feynman diagram splitting. I also cover some more recent optimizations of SIMD color sums. In Sec. 4, I describe the various options I introduced to split the calculation of Feynman diagrams into different groups of diagrams and/or different kernels, and different source code files. I also give the first CUDACPP performance results for physics processes with large numbers of final state gluons, which could be computed thanks to these new features. In Sec. 5, I finally give my conclusions and an an outlook for this work.
@@ 2.1 Three software areas and development phases
    ME engines: the port to GPUs and SIMD CPUs, using CUDA/HIP and vectorized C++, of the computational engines to calculate MEs from particle momenta, and their backport to the code-generating Python framework, for different physics processes.
@@
More specifically, in all of the work that I present in this paper, I only developed code using the standalone application check.exe, where all three main components of an event generation application (random numbers, phase space sampling and ME calculation) are implemented in C++ and CUDA using a data parallel approach. This is because check.exe makes it possible to focus on the optimization of the ME engine without being constrained by the MadEvent Fortran infrastructure, where the legacy phase space sampling and other sequential non-ME components slow down the whole workflow. I routinely built check.exe using the CUDACPP makefile cudacpp.mk, without even building MadEvent or other Fortran code (which proceeds via a separate makefile that internally delegates the build of the CUDACPP library to cudacpp.mk). I did however validate my changes by running the full MadEvent test suite at the end of each development cycle.
@@
Table 1: Pseudo-code of the $gg\!\rightarrow\!t\bar{t}$ ME calculation in CUDACPP before any kernel splitting changes (“ihel0”). Function main represents the check.exe standalone application. Note that sigmaKin is the only GPU kernel (__global__ function) in this workflow. Only a simplified GPU version of the code is shown, using scalar fptype floating point and cxtype complex number types: the actual code that supports both GPUs and SIMD CPUs uses fptype_sv and cxtype_sv data types inside calculate_wavefunction, making it possible to use the same formal code for scalar data and for SIMD vectors (see Ref. bib:vchep2021 for details). The CPU branch of the code for main, encapsulated by an #ifdef and not shown here, uses the Host versions of various classes instead of the Device versions. The pseudo-code for the color sum schematically indicates that this proceeds via a quadratic form using a color matrix, and that color amplitudes are converted from fptype/cxtype floating point precision to a potentially different fptype2/cxtype2 precision (this enables the “mixed” precision mode, where the former is based on double and the latter on float).